OpenStudy (anonymous):

Write the equation of a line parallel to the line 2x + 3y = 5 and passes through the point (9, -3)

OpenStudy (anonymous):

A. y equals negative two-thirds x plus 3 B. y equals negative two-thirds x plus five-thirds C. y = 2x – 21 D. y equals negative two-thirds x minus 9

OpenStudy (anonymous):

2x + 3y = 5 -parallel=same slope but different y intercept so 2x+3y=5 3y=-2x+5 y=(-2/3)x+5/3 slope= -2/3\ y=(-2/3)x+b and we have point (9, -3) where x=9 and y=-3 -3=(-2/3)(9)+b -3=-18/3+b -3=-6+b 3=b so the equation is y=(-2/3)x+3

OpenStudy (anonymous):

so A